For any positive intege rx, let f(x) =x^x. Suppose that n is a positive integer such that there exists a positive integer m with m not equal 1 such that f(f(f(m))) =m^m^n+2020. Compute the smallest possible value of n.
1
Expert's answer
2020-12-15T01:20:29-0500

"f(f(f(m))) = f(f(m^m))=f((m^m)^{m^m})=f(m^{m^{m+1}})=(m^{m^{m+1}})^{(m^{m^{m+1}})}=m^{m^{1+m^{m+2}}}"


Then find n by substituting it in equation:


"f(f(f(m)))=m^{m^n}+2020 \\
m^{m^{1+m^{m+2}}}=m^{m^n}+2020"

"n = log_m{[log_m{(m^{m^{1+m^{m+2}}}-2020)}]}" -- is the strongly increased function and

"m^{m+2}<n<1+m^{m+2}"


Hence, the smallest possible value of n could be when m = 2: 8<=n<=9 and then "m^n\ge2^8>2"


But from other hand

"m^{m^{1+m^{m+2}}}=m^{m^n}+2020 \\
m^{m^{1+m^{m+2}}}-m^{m^n}=2^2*5*101\\
m^{m^n}*(m^{m^n(m^{1+m^{m+2}-n}-1)}-1)=2^2*5*101"

The left side number has prime factors with power more than 2: "m^n>2^8>2" - contradiction.


Hence, n is not exist